OpenStudy (skyepurple1997):

Based on data from a statistical abstract, only about 10% of senior citizens (65 years old or older) get the flu each year. However, about 23% of the people under 65 years old get the flu each year. In the general population, there are 13.5% senior citizens (65 years old or older). (Round your answers to three decimal places.) (a) What is the probability that a person selected at random from the general population is senior citizen who will get the flu this season?

OpenStudy (skyepurple1997):

(b) What is the probability that a person selected at random from the general population is a person under age 65 who will get the flu this year? (c) Repeat parts (a) and (b) for a community that has 85% senior citizens. (d) Repeat parts (a) and (b) for a community that has 45% senior citizens

OpenStudy (skyepurple1997):

@phi

OpenStudy (phi):

To keep the numbers simple, let's say there are 100 people (and we allow "fractions" of a person ) From ** In the general population, there are 13.5% senior citizens (65 years old or older)*** we divide the population into 86.5 young 13.5 old (note this adds up to 100) ***10% of senior citizens (65 years old or older) get the flu each year. *** so 10% of 13.5 get the shot: 1.35 and the rest do not get a shot: 90% of 13.5 = .9 * 13.5 = 12.15 *** 23% of the people under 65 years old get the flu each year*** that means 0.23 * 86.5 = 19.9 young get a shot and .77 *86.5 = 66.6 young do not get a shot

OpenStudy (phi):

now form a table showing the different categories young old shot 19.9 1.35 no shot 66.6 12.15 **** (a) What is the probability that a person selected at random from the general population is senior citizen who will get the flu this season?**** probability is "number of outcomes that meet the criteria" divided by total number the number of people in the old/shot category is 1.35 total number is 100. so probability is 1.35/100 or 1.35% chance of picking an old person who got a shot.